### v1

# Aligning Large Language Models for Controllable Recommendations

Abstract
Inspired by the exceptional general intelligence of Large Language Models (LLMs), researchers have begun to explore their application in pioneering the next generation of recommender systems — systems that are conversational, explainable, and controllable. However, existing literature primarily concentrates on integrating domain-specific knowledge into LLMs to enhance accuracy, often neglecting the ability to follow instructions. To address this gap, we initially introduce a collection of supervised learning tasks, augmented with labels derived from a conventional recommender model, aimed at explicitly improving LLMs’ proficiency in adhering to recommendation-specific instructions. Subsequently, we develop a reinforcement learning-based alignment procedure to further strengthen LLMs’ aptitude in responding to users’ intentions and mitigating formatting errors. Through extensive experiments on two real-world datasets, our method markedly advances the capability of LLMs to comply with instructions within recommender systems, while sustaining a high level of accuracy performance.
Aligning Large Language Models for Controllable Recommendations
[…]
1 Introduction
Recommender systems are designed to identify and suggest the most appropriate items to users from a vast array of candidates, based on the users’ profiles, past interactions, and present intentions. Witnessing the impressive capabilities of Large Language Models (LLMs), such as knowledge retention, reasoning, and problem-solving, researchers are now exploring the integration of LLMs into the next wave of intelligent recommender systems, which aim to be conversational, explainable, and controllable. Bridging the gap between the general-purpose LLMs and the specific requirements of recommendation tasks poses a challenge. In this context, fine-tuning LLMs with domain-specific knowledge and recommendation-focused tasks emerges as a promising strategy to harness their potential for advanced recommendation purposes Bao et al. (2023); Zhang et al. (2023); Chen (2023).
Typical approaches in literature involve reformatting recommendation tasks — such as item reranking or click-through rate (CTR) prediction — into natural language constructs to facilitate the fine-tuning of LLMs. However, we have observed that LLMs fine-tuned through this straightforward method, albeit enhancing accuracy in offline evaluations, frequently generate outputs with domain-specific formatting errors. These errors may manifest as repeated items in the top-k recommendations or the inclusion of items previously interacted with by the user. Additionally, these LLMs exhibit a limited ability to adhere to user instructions. This compromises their effectiveness as viable agents in real-world recommender systems. A vidid example can be found in Section 3.4.
In this paper, we investigate the alignment of an LLM for recommender systems. Our objective extends beyond merely improving the recommendation accuracy of an original LLM; we aim to significantly enhance controllability and reduce formatting errors. Drawing inspiration from the Reinforcement Learning from Human Feedback (RLHF) framework Ouyang et al. (2022), our methodology is structured into two phases: the supervised learning (SL) stage and the reinforcement learning (RL) stage. To inject domain-specific knowledge and foster recommendation-relevant control capabilities within the LLM, we devise a series of fine-tuning tasks, including item recommendation, item search, category control, and category proportion control. These tasks often necessitate generating a list of items that not only meet user instructions but also maintain high quality, despite the typically sparse ground-truth signals found in user behavior history. To tackle this issue, we propose augmenting supervised labels using predictions from traditional recommender models, such as SASRec Kang and McAuley (2018).
After the SL stage, the LLM exhibits a significantly enhanced ability to follow recommendation-related instructions, surpassing existing approaches that solely fine-tune the LLM on item recommendation and search tasks. Nevertheless, the SL stage’s data generation process inherently provides only positive examples for each instruction. To address scenarios where the LLM generates suboptimal responses, we introduce an RL stage with carefully crafted reward signals to further refine the LLM’s capacity to follow instructions. To our knowledge, this is the first study that employs both SL and RL stages to align LLMs for recommendation purposes. We conduct comprehensive experiments on two real-world datasets, Steam and Amazon Movie, with results demonstrating that our method markedly improves the LLM’s ability to follow instructions while simultaneously reducing formatting errors. Our major contributions are summarized as follows:
  • We introduce a novel supervised learning stage, which encompasses a suite of tasks designed for enhancing controllability and label-augmentation by a teacher recommender model, to align an LLM into an intelligent recommender agent.
  • To mitigate formatting errors and improve the instruction-following ability, We further design an alignment stage based on reinforcement learning with a variety of rewards that are tailored for the nuances of the recommendation task.
  • Experiment results validate that our model markedly surpasses existing LLM-based recommendation models, exhibits a robust capacity to follow users’ instructions while maintaining a high level of recommendation precision.
2 Methodology
2.1 Intention Categories
This paper aims to enhance the instruction-following capabilities of LLMs for recommendation tasks. We categorize recommendation instructions into three distinct types:
Implicit intention. This is the assumed default where the prompt describes the user’s profile (such as attributes and past favored items). The LLM is tasked with recommending items that align with the user’s preferences.
Item-wise intention. In addition to the user profile, users may express specific desires, such that the recommended items should either exhibit particular characteristics (“I wish to watch an action movie”) or exclude them (“Please avoid suggesting any action movies”).
List-wise intention. Users may have requirements for the entire list of recommended items; hence, evaluating individual items’ attributes is insufficient. For example, if all items in a recommendation list belong to the same category A, the user may be disappointed by the lack of diversity. Consequently, the user might request the recommender system to ensure that the proportion of category A is below a certain threshold.
To effectively train LLMs as recommender agents capable of adhering to these three types of instructions, we introduce a novel two-stage fine-tuning approach: a supervised learning (SL) stage (Section2.2) followed by a reinforcement learning (RL) stage (Section2.3). The overall framework is illustrated in Figure 1.
Figure 1: Overview of the proposed method.
2.2 The SL Stage
We represent each data sample in natural language, adopting the format “Instruction: [Prompt_Content]. Output: [Response_Content]”, where [Prompt_Content] includes the detailed instruction reflecting the user’s profile and intention, and [Response_Content] contains the recommendations that fulfill the instruction. Different from traditional recommender systems that utilize item IDs, we employ item titles in both [Prompt_Content] and [Response_Content] to ensure comprehension by the LLM. Data samples are generated according to the following tasks:
2.2.1 Data Generation Tasks
Sequential Recommendation Instructions ($\bm{I_{0}}$)
This task represents the traditional sequential recommendations: given a user’s previously interacted items, the goal is to predict future interactions. Specifically, we use the first $n-1$ items to construct the user’s behavioral profile, while the $n^{th}$ item serves as the ground-truth label. The LLM is instructed to recommend $k$ items; if the ground-truth item is among these $k$ suggestions, the recommendation is considered a successful hit. The value of $k$ for each data sample is randomly selected from $1$ to $10$.
Category Control Instructions ($\bm{I_{1}}$)
This task corresponds to instructions with item-wise intention, which we implement into two distinct types: 1) Positive control ($\bm{I_{1}^{+C}}$), where the user hopes to receive more items in the recommendation list that match the specific category $C_{target}$. 2) Negative control ($\bm{I_{1}^{-C}}$), where users indicate a weariness towards a certain category $C_{target}$ and wish to minimize the inclusion of such items in their received recommendations as much as possible.
Category Proportion Control Instructions ($\bm{I_{2}}$)
We implement the list-wise intention into three distinct types: 1) $\bm{I_{2}^{CP\leq*}}$, in this case, the user hopes to have the proportion of the item of $C_{target}$ less than a certain value. 2) $\bm{I_{2}^{CP\approx*}}$, in this case, the user hopes that the proportion of items in $C_{target}$ will be approximately a certain value. 3) $\bm{I_{2}^{CP\geq*}}$, in this case, the user hopes that the proportion of items in $C_{target}$ is greater than a certain value.
Item Search Instructions($\bm{I_{3}}$)
To aid the LLM in memorizing item attributes, we introduce an item search task: the objective is to retrieve $k$ items belonging to a target category $C_{target}$. For this purpose, we randomly select $k$ items from $C_{target}$ to serve as the ground truth for the response.
ShareGPT($\bm{I_{4}}$)
To avoid catastrophic forgetting and preserve the general intelligence capabilities of the LLM, we follow Zeng et al. (2023) and integrate a specific proportion of ShareGPT training data into the SL stage.
2.2.2 Label Augmentation
Instructions $\bm{I_{0}}$, $\bm{I_{1}}$, and $\bm{I_{2}}$ require the inclusion of $k$ items within the response text. However, due to the typically sparse nature of user historical behavior, it is often impractical to construct a ground-truth response based solely on this information. To overcome this limitation, we employ the sequential recommender model SASRec Kang and McAuley (2018) as a teacher model to generate a set of top recommendations, $\bm{P}_{SASRec}$, for each data sample. We then curate the top-$k$ list by selecting items from $\bm{P}_{SASRec}$ that align with the given instructions. The top-$k$ list is assembled as follows: the first item is the ground-truth item (i.e., the $n^{th}$ item in the user’s history). The method for compiling the remaining $k-1$ items varies slightly: for $\bm{I_{0}}$, we fill the list with SASRec’s top predictions; for $\bm{I_{1}}$ and $\bm{I_{2}}$, we filter SASRec’s predictions to ensure the final recommendation list adheres to the specified instruction.
2.3 The RL Stage
Following previous works Touvron et al. (2023); Ouyang et al. (2022), we employ the Proximal Policy Optimization (PPO) algorithm Schulman et al. (2017) to further fine-tune the model after the SL stage. Different from Ouyang et al. (2022), scores are not produced by a reward model; instead, they are derived from reward rules that are specifically tailored for $\bm{I_{0}}$, $\bm{I_{1}}$, and $\bm{I_{2}}$. Fundamentally, rewards consist of two components: item-level rewards and list-level rewards.
2.3.1 Item-level Reward
The item-level reward assigns a score to each item generated by the LLM, serving as an immediate reward of reinforcement.
For each $item_{i}$ in LLM’s recommendation list, let $Rank_{i}$ refer to the rank of $item_{i}$ in the teacher model (SASRec)’s prediction list. We can then calculate the preference scores for $item_{i}$ by:
$Scores_{i}=\begin{cases}-1,&\text{if $item_{i}$ is illegal}\\ +1,&\text{if $item_{i}$ is $item_{target}$}\\ \frac{1}{log_{2}(Rank_{i}+3)},&\text{else}\end{cases}$
$item_{i}$ is considered illegal if it meets any of the following conditions: $item_{i}$ does not exist, it is a duplicate of any item within the preceding set $item_{[1,...,i-1]}$, it is identical to an item in the user’s history, or its index $i$ exceeds $k$.
Beyond the preference score, $Scores$, a control effect score, $Scores^{ctl}$, is required to gauge the extent to which an item corresponds with user intentions. Essentially, a generated item that adheres to the given instruction is awarded a high score, signifying positive reinforcement, while a non-conforming item incurs a low score, serving as negative feedback. The complete calculation is described in algorithm 1 in the Appendix. Finally, we get item-level reward $R_{item}$, which measures the overall merits of each item:
$R_{item}=0.5*Scores+0.5*Scores^{ctl}$ (2)
2.3.2 List-level Reward
The list-wise reward, which measures how well the entire list of recommendations matches the user’s preferences and control intentions, is added to the last token of output, serving as a termination reward. To encourage the LLM to rank ground-truth items in top positions, we adjust the $Scores$ as per Equation 3 to derive $Score^{*}$:
$Scores_{i}^{*}=\begin{cases}-1,&\text{if $item_{i}$ is illegal}\\ \frac{Scores_{i}}{log_{2}(i+2)},&\text{else}\end{cases}$ (3)
In addition to this, we need to measure how well the output matches the control intention. Let $Count_{in},Count_{out}$ denote the number of items that belong/do not belong to the target category. For different control intentions, we use different calculation methods to get $Score^{ctl}_{list}$:
$Score^{ctl}_{list}=\begin{cases}\mathrm{sum}(Scores^{*}),&\text{if $I_{0}$}\\ \frac{1}{log_{2}((k-Count_{in})+2)},&\text{if $I_{1}^{+C}$}\\ \frac{1}{log_{2}((k-Count_{out})+2)},&\text{if $I_{1}^{-C}$}\\ \frac{1}{log_{2}(max(Count_{in}-k*m,0)+2)},&\text{if $I_{2}^{CP\leq m}$}\\ \frac{1}{log_{2}(max(k*m-Count_{in},0)+2)},&\text{if $I_{2}^{CP\geq m}$}\\ \frac{1}{log_{2}(abs(Count_{in}-k*m)+2)},&\text{if $I_{2}^{CP\approx m}$}\\ \end{cases}$
Finally, we get $R_{list}$, which measures the overall merits of the recommendation list:
$R_{list}=0.5*\mathrm{sum}(Scores^{*})+0.5*Score^{ctl}_{list}$
3 Experiments
3.1 Experiment Setting
3.1.1 Dataset
We experiment with two datasets in recommender system domain: the Amazon Movies & TVs ***https://cseweb.ucsd.edu/~jmcauley/datasets/amazon˙v2/ dataset and the Steam dataset Kang and McAuley (2018). Both datasets include item category information, which aids in constructing the user’s control intentions. The statistics of our dataset are summarized in Table 1. We employed the leave-one-out approach to split the dataset. Therefore, the size of test set is consistent with the number of users in each dataset. To accelerate the validation process, we only include four types of instructions in the valid set: $I_{0}$, $I_{1}^{+C}$, $I_{1}^{-C}$, and $I_{2}^{CP\approx 50\%}$. Each type of instruction has 320 instances. Upon observing multiple metrics on the validation set, we find that SL typically converged around 30 epochs. Full details of training set on each instruction task can be found at Table 7 in the Appendix.
$Dataset$$\mathrm{\#Users}$$\mathrm{\#Items}$$\mathrm{\#Inters}$$\mathrm{\#Sparsity}$
$Movie$ $13218$ $18744$ $744313$ $99.70\%$
$Steam$ $12658$ $8572$ $632900$ $99.42\%$
Table 1: General Statistics of the Two Datasets
3.1.2 Implementation Details
We choose Llama-2-7b-chat as the foundational model for our research Touvron et al. (2023). We set model’s maximum sequence length to 1024 tokens. User behavior sequences are truncated to incorporate no more than 10 items, and item titles that are excessively lengthy are condensed to a maximum of 64 tokens. Furthermore, to accommodate the complete recommendation list within the output, we reserve a larger token count, setting the maximum output length of our model to 512 tokens. Instructions and responses are formatted using the official prompt template in Touvron et al. (2023).
We use LoRA Hu et al. (2022) to fine-tune all linear layers in Llama2-7b-chat, with trainable parameters accounting for about 0.6% of the total parameters in SL and 0.3% in RL. The optimizer is Adam. In the SL stage, the learning rate is set to 0.001, LoRA dimmention is set to 16, LoRA alpha to 8, and batch size is 64. The ShareGPT data accounted for 50% of the total training data. In the RL stage, the learning rate is $5\times 10^{-6}$. We set separate LoRA layers for actor and critic networks, with LoRA dimension at $4$ and LoRA alpha at $2$. To encourage model exploration, we set the temperature to 0.7 and the weight of entropy loss to 0.01. To ensure that the information in $R_{list}$ is not overshadowed by $R_{item}$, we amplify $R_{list}$ by a factor of ten. Source code will be released at https://github.com/microsoft/recai.
3.1.3 Metrics
For all models, we use Top-k hit ratio (HR@K) and Top-k NDCG (NDCG@K) to evaluate the accuracy of recommendations. Meanwhile, we use some additional indicators to evaluate the model’s ability to follow user instructions: 1). For category control, we use the Top-k target category proportion (TCP@K) metric to evaluate the proportion of target categories in the recommendation list. The calculation of TCP@K is shown in Equation 6. 2). For category proportion control, we use the target category proportion accuracy (CPA), defined in Equation 7, to measure whether the responding item distribution complies with the instruction. Note that $k$ indicates the number of recommended items and $m$ indicates the proportion requirement in instructions.
$TCP@K=\frac{1}{K}\sum_{i=1}^{K}{\mathds{1}(item_{i}\in C_{target})}$ (6)
$CPA=\begin{cases}\mathds{1}(Count_{in}\leq k*m),&\text{if $I_{2}^{CP\leq m}$}\\ \mathds{1}(Count_{in}\geq k*m),&\text{if $I_{2}^{CP\geq m}$}\\ \mathds{1}(\mathrm{abs}(Count_{in}-k*m)\leq 1),&\text{if $I_{2}^{CP\approx m}$}\\ \end{cases}$
3.1.4 Baselines
We divide the compared methods into 3 classes: general LLM, fine-tuned LLM, and our method and its variants. For sequential recommendation, we additionally compared with SASRec Kang and McAuley (2018) that we used as teacher model.
General LLM:
We use a closed-source LLM GPT-3.5-turbo-0613 https://platform.openai.com/docs/models/gpt-3-5-turbo (GPT-3.5 for short) and an open-source LLM Llama2-7b-chat Touvron et al. (2023). Due to the high cost, all GPT-3.5 results are obtained from the first $1{,}000$ samples of the test set.
Fine-tuned LLM:
InstructRec Zhang et al. (2023) uses $39$ types of recommendation-related instructions constructed from user history and review data. It uses Beam search to obtain the top-k recommendation list during the test phase. Its backbone is 3B Flan-T5-XL. PALR Chen (2023) also uses Llama2-7b-chat as the backbone. We fine-tune both models using our datasets, employing data sample generation techniques as described in their respective papers. For instance, a key distinction between PALR and our supervised instruction tuning task, $I_{\{0\}}$, is that PALR does not utilize a teacher recommender model for label augmentation; instead, it relies solely on user behavior history to derive label responses.
Our method and variants:
$\mathbf{Ours_{v1}}$ is a model that only uses sequential recommendation instructions ($I_{\{0\}}$) in SL. $\mathbf{Ours_{v2}}$ is a model uses all instructions ($I_{\{0,1,2,3,4\}}$) in SL. $\mathbf{Ours_{v3}}$ is a model obtained by further fine-tuning $\mathrm{Ours_{v2}}$ using RL, but without the item-level reward. $\mathbf{Ours_{full}}$ is the complete version of our method. We compare different variants of our method as an ablation study to verify the effect of different components.
3.2 Overall Performance
3.2.1 Sequential Recommendation
We begin by assessing model performance under standard sequential recommendation scenarios, where users have no explicit intentions. It is important to clarify that our objective in this paper is to improve the ability of LLMs to follow recommendation-related instructions, not to outperform the teacher model, SASRec. Achieving recommendation accuracy on par with the teacher model is considered satisfactory for our purposes. Table 2 presents the comprehensive results. Initially, all fine-tuned models surpass the performance of general LLMs such as GPT-3.5 and Llama2-7b, affirming the necessity of fine-tuning for domain-specific tasks. Additionally, our approach significantly outperforms the fine-tuned benchmarks (InstructRec and PALR), validating the efficacy of our Supervised Learning (SL) stage. Lastly, our completed model, $\mathbf{Ours_{full}}$, achieves results comparable to the teacher model, SASRec, reinforcing the premise that an LLM must first grasp user preferences before its instruction-following capabilities can be further evaluated.
Dataset Method $\mathrm{HR@10}$ $\mathrm{NDCG@10}$ $\mathrm{HR@5}$ $\mathrm{NDCG@5}$
Movie $\mathrm{SASRec}$ 0.1229 $\underline{0.0913}$ $\underline{0.1018}$ $\underline{0.0844}$
$\mathrm{GPT-3.5}$ $0.0050$ $0.0025$ $0.0030$ $0.0019$
[…]
$\mathrm{InstructRec}$ $0.0524$ $0.0381$ $0.0406$ $0.0343$
$\mathrm{PALR}$ $0.0868$ $0.0787$ $0.0832$ $0.0775$
$\mathrm{Ours_{v1}}$ $0.1108$ $0.0861$ $0.0991$ $0.0827$
$\mathrm{Ours_{v2}}$ $\underline{0.1211}$ 0.0927 0.1056 0.0880
$\mathrm{Ours_{v3}}$ $0.1150$ $0.0858$ $0.1010$ $0.0824$
$\mathrm{Ours_{full}}$ $0.1148$ $0.0867$ $0.1001$ $0.0825$
Steam $\mathrm{SASRec}$ 0.1121 0.0648 0.0778 0.0538
$\mathrm{GPT-3.5}$ $0.0160$ $0.0079$ $0.0090$ $0.0055$
[…]
$\mathrm{InstructRec}$ $0.0220$ $0.0113$ $0.0122$ $0.0082$
$\mathrm{PALR}$ $0.0408$ $0.0320$ $0.0373$ $0.0308$
$\mathrm{Ours_{v1}}$ $0.0930$ $0.0535$ $0.0666$ $0.0451$
$\mathrm{Ours_{v2}}$ $\underline{0.1036}$ $\underline{0.0583}$ $\underline{0.0717}$ $\underline{0.0479}$
$\mathrm{Ours_{v3}}$ $0.1014$ $0.0557$ $0.0695$ $0.0454$
$\mathrm{Ours_{full}}$ $0.1001$ $0.0551$ $0.0688$ $0.0449$
Table 2: Results of ($I_{0}$). The best result is highlighted in boldface and the runner-up is denoted with underline.
3.2.2 Category Control
To evaluate the Category Control Instructions ($\bm{I_{1}}$), we use the following two settings:
Positive ($I_{1}^{+C}$): We utilize the target item’s category as the control signal $C_{target}$ and incorporate a positive descriptor of $C_{target}$ into the input instruction to represent the user’s explicit intent to receive more items within that category.
Negative ($I_{1}^{-C}$): We analyze the output statistics of top-$k$ recommendations by SASRec to identify the category with the largest proportion — excluding the target item’s category — as $C_{target}$ for simulated control. A negative descriptor of $C_{target}$ is then embedded in the instruction to reflect the user’s intent to minimize items from this category.
$\bm{Control}$ $\bm{I_{1}^{+C}}$ $\bm{I_{1}^{-C}}$
Dataset Model $\mathrm{HR@10}$ $\mathrm{NDCG@10}$ $\mathrm{TCP@10(\%)}\uparrow$ $\mathrm{HR@10}$ $\mathrm{NDCG@10}$ $\mathrm{TCP@10(\%)}\downarrow$
Movie $\mathrm{GPT-3.5}$ $0.0200$ $0.0111$ $7.39(2.23)$ $0.0150$ $0.0074$ $7.78(19.72)$
[…]
$\mathrm{InstructRec}$ $0.1045$ $0.0687$ $37.39(7.16)$ $0.0362$ $0.0242$ $56.61(41.82)$
$\mathrm{PALR}$ $0.0832$ $0.0746$ $8.70(7.64)$ $0.0809$ $0.0728$ $46.64(61.46)$
$\mathrm{Ours_{v1}}$ $0.1054$ $0.0788$ $8.69(8.07)$ $\underline{0.1023}$ $\underline{0.0778}$ $35.99(40.83)$
$\mathrm{Ours_{v2}}$ 0.2620 0.1814 $81.38(8.05)$ 0.1099 0.0832 $\underline{19.35}(39.80)$
$\mathrm{Ours_{v3}}$ $0.2383$ $0.1609$ $\underline{89.06}(7.89)$ $0.1017$ $0.0772$ $15.80(39.04)$
$\mathrm{Ours_{full}}$ $\underline{0.2336}$ $\underline{0.1574}$ $\textbf{93.52}(7.81)$ $0.0999$ $0.0756$ $\textbf{11.49}(38.41)$
Steam $\mathrm{GPT-3.5}$ $0.0360$ $0.0187$ $37.62(25.59)$ $0.0090$ $0.0044$ $19.67(43.59)$
[…]
$\mathrm{InstructRec}$ $0.0494$ $0.0247$ $51.70(29.47)$ $0.0114$ $0.0059$ $70.39(53.49)$
$\mathrm{PALR}$ $0.0392$ $0.0301$ $13.04(12.45)$ $0.0350$ $0.0275$ $47.01(52.60)$
$\mathrm{Ours_{v1}}$ $0.0922$ $0.0509$ $27.23(26.41)$ $0.0887$ $0.0488$ $51.38(56.63)$
$\mathrm{Ours_{v2}}$ 0.3484 0.2110 $92.25(28.02)$ 0.1236 0.0724 $12.18(55.67)$
$\mathrm{Ours_{v3}}$ $\underline{0.3422}$ $\underline{0.2049}$ $\underline{95.13}(29.75)$ $\underline{0.1200}$ $\underline{0.0696}$ $\underline{6.42}(54.37)$
$\mathrm{Ours_{full}}$ $0.3395$ $0.2036$ $\textbf{95.80}(29.92)$ $0.1167$ $0.0676$ $\textbf{5.51}(54.00)$
Table 3: Results of $I_{1}$ control. The best result is highlighted in boldface and the runner-up is denoted with underline. Values in parentheses show outcomes of the corresponding model without control signals.
Table 3 presents the results, using the TCP@10 metric to assess conformity to control signals, while HR@10 and NDCG@10 indicate recommendation accuracy when explicit intentions are additionally included in the prompt. Values in parentheses show the outcomes of each corresponding model without control signals. Two key observations emerge: Firstly, differentiating supervised learning tasks by intention is essential. Evidence for this includes InstructRec’s improved TCP performance in the $I_{1}{+C}$ task, which aligns with its training, versus no effect in the $I_{1}{-C}$ task not covered by its training. Similarly, $\mathbf{Ours_{v1}}$, trained solely on recommendation tasks, underperforms in TCP compared to $\mathbf{Ours_{v2}}$. Secondly, $\mathbf{Ours_{v2}}$, $\mathbf{Ours_{v3}}$, and $\mathbf{Ours_{full}}$ show notable TCP enhancements, with $\mathbf{Ours_{full}}$ leading. Additionally, HR and NDCG metrics also see a significant lift compared to Table 2, confirming our method’s effectiveness in adhering to category control instructions.
3.2.3 Category Proportion Control
To evaluate the Category Control Instructions ($\bm{I_{1}}$), we use the following three settings:
[…]
Results are reported in Table 4. In this case, the primary metric is CPA. Overall, $Ours_{full}$ achieves the best performance, outperforming all other baselines and its own variants.
[…]
Dataset Model $\mathrm{HR@10}$ $\mathrm{NDCG@10}$ $\mathrm{CPA(\%)}\uparrow$ $\mathrm{HR@10}$ $\mathrm{NDCG@10}$ $\mathrm{CPA(\%)}\uparrow$ $\mathrm{HR@10}$ $\mathrm{NDCG@10}$ $\mathrm{CPA(\%)}\uparrow$
Movie $\mathrm{GPT-3.5}$ $0.0050$ $0.0021$ $1.80(0.50)$ $0.0210$ $0.0124$ $9.71(0.70)$ $0.0250$ $0.0167$ $6.61(1.10)$
[…]
$\mathrm{InstructRec}$ $0.0372$ $0.0252$ $9.75(15.87)$ $0.1031$ $0.0684$ $31.24(3.96)$ $0.1041$ $0.0692$ $\underline{57.58}(9.60)$
$\mathrm{PALR}$ $0.0792$ $0.0711$ $29.76(6.70)$ $0.0813$ $0.0724$ $10.28(3.16)$ $0.0825$ $0.0733$ $11.17(9.84)$
$\mathrm{Ours_{v1}}$ $\underline{0.1018}$ 0.0742 $30.97(14.34)$ $0.1048$ $0.0766$ $16.14(4.43)$ $0.1061$ $0.0776$ $10.46(10.19)$
$\mathrm{Ours_{v2}}$ 0.1027 $\underline{0.0688}$ $29.18(14.93)$ 0.2034 0.1417 $\underline{58.71}(4.61)$ $\underline{0.2055}$ $\underline{0.1433}$ $45.89(9.96)$
$\mathrm{Ours_{v3}}$ $0.0943$ $0.0594$ $\underline{33.74}(15.03)$ $0.1948$ $0.1319$ $65.78(4.58)$ $0.1994$ $0.1340$ $48.19(9.21)$
$\mathrm{Ours_{full}}$ $0.0981$ $0.0642$ $\textbf{45.90}(16.33)$ $\underline{0.2005}$ $\underline{0.1354}$ $\textbf{69.81}(4.68)$ 0.2134 0.1447 $\textbf{61.98}(9.15)$
Steam $\mathrm{GPT-3.5}$ $0.0170$ $0.0074$ $29.80(14.40)$ $0.0290$ $0.0152$ $25.60(3.40)$ $0.0360$ $0.0202$ $51.10(39.60)$
[…]
$\mathrm{InstructRec}$ $0.0109$ $0.0056$ $15.86(21.39)$ $0.0442$ $0.0223$ $13.22(3.15)$ $0.0435$ $0.0220$ $55.90(40.30)$
$\mathrm{PALR}$ $0.0333$ $0.0257$ $6.18(2.47)$ $0.0374$ $0.0287$ $13.39(4.13)$ $0.0388$ $0.0296$ $20.11(18.96)$
$\mathrm{Ours_{v1}}$ $0.0863$ $0.0478$ $20.51(12.25)$ $0.0897$ $0.0496$ $12.23(3.96)$ $0.0903$ $0.0502$ $39.22(38.73)$
$\mathrm{Ours_{v2}}$ $0.1144$ $0.0647$ $41.26(14.90)$ $0.2072$ $0.1191$ $71.97(3.26)$ $0.2184$ $0.1238$ $70.79(39.79)$
$\mathrm{Ours_{v3}}$ $\underline{0.1172}$ $\underline{0.0663}$ $\underline{65.41}(18.04)$ $\underline{0.2105}$ $\underline{0.1206}$ $\textbf{77.51}(2.14)$ $\underline{0.2283}$ $\underline{0.1286}$ $\underline{82.07}(40.69)$
$\mathrm{Ours_{full}}$ 0.1183 0.0663 $\textbf{70.58}(18.15)$ 0.2225 0.1295 $\underline{74.29}(2.77)$ 0.2382 0.1365 $\textbf{88.40}(40.98)$
Table 4: Results of $I_{2}$ control. The best result is highlighted in boldface and the runner-up is denoted with underline. Values in parentheses show outcomes of the corresponding model without control signals.
3.3 Formatting and General Evaluation
Finally, we assess the formatting ability and overall linguistic capabilities of LLMs in generating structured recommendations. We measure formatting ability for the recommendation domain in four dimensions: 1) CorrectCount: The accuracy of the recommended item count against the given $k$ in the instruction. 2) RepeatItem: The frequency of repeated items in the recommendation list. 3) NonExist: The occurrence of non-existent (hallucinated) items in the list. 4) InHistory: The rate of recommended items already present in the user’s history. For a more challenging test, LLMs are tasked with recommending $k$ items where $k$ is a random number between 11 and 15, despite training only on ranges from 1 to 10. To evaluate the general language ability of LLMs, we utilize two standard tasks, MMLU with five examples (5-shot) and GSM8K with eight examples (8-shot), leveraging the unified framework provided by https://github.com/EleutherAI/lm-evaluation-harness.
Dataset Method Formatting Quality(%) Precision Generalization
[…]
Movie $\mathrm{GPT-3.5}$ $100.00$ $2.45$ $66.10$ $4.22$ $0.0100$ $0.0034$ $0.700$ $0.7460$
[…]
$\mathrm{InstructRec}$ $100.00$ $10.01$ $8.52$ $15.35$ $0.0546$ $0.0378$- --
$\mathrm{PALR}$ $77.27$ $65.92$ $4.06$ $9.31$ $0.0869$ $0.0787$ $0.377$ $0.1099$
$\mathrm{Ours_{v1}}$ $92.96$ $13.63$ $7.03$ $5.76$ $0.1164$ $0.0876$ $0.341$ $0.1318$
$\mathrm{Ours_{v2}}$ $100.00$ $9.61$ $5.27$ $4.02$ 0.1285 0.0950 $0.450$ $\underline{0.1842}$
$\mathrm{Ours_{v3}}$ $\underline{100.00}$ $\underline{2.37}$ $\underline{1.14}$ $\underline{1.36}$ $0.1214$ $0.0886$ $\underline{0.453}$ $0.1789$
$\mathrm{Ours_{full}}$ 100.00 1.14 0.95 1.24 $\underline{0.1220}$ $\underline{0.0890}$ 0.455 $0.1782$
Steam $\mathrm{GPT-3.5}$ $99.90$ $2.44$ $26.76$ $4.50$ $0.0200$ $0.0094$ $0.700$ $0.7460$
[…]
$\mathrm{InstructRec}$ $98.41$ $0.99$ $4.60$ $7.54$ $0.0270$ $0.0130$- --
$\mathrm{PALR}$ $97.53$ $17.67$ $46.78$ $2.88$ $0.0404$ $0.0316$ $0.417$ $0.1327$
$\mathrm{Ours_{v1}}$ $95.79$ $3.95$ $3.00$ $1.49$ $0.1029$ $0.0559$ $0.327$ $0.0819$
$\mathrm{Ours_{v2}}$ $100.00$ $2.68$ $1.59$ $2.55$ 0.1152 0.0612 0.458 $\underline{0.2146}$
$\mathrm{Ours_{v3}}$ $\underline{100.00}$ $\underline{0.37}$ $\underline{1.04}$ $\underline{0.22}$ $0.1149$ $\underline{0.0593}$ $0.457$ $0.2039$
$\mathrm{Ours_{full}}$ 100.00 0.23 0.78 0.17 $\underline{0.1149}$ $0.0589$ $\underline{0.457}$ $0.2123$
Table 5: Results of formatting and general evaluation. The best result (excluding GPT-3.5) is highlighted in boldface and the runner-up is denoted with underline.
Table 5 presents the comprehensive outcomes. Notably, nearly all models, with the exception of PALR on the Movie dataset, excel in the CorrectCount metric. For other formatting metrics such as RepeatItem, NonExist, and InHistory, a progressive enhancement is evident from $Ours_{v1}$ through $Ours_{v3}$, culminating in $Ours_{full}$. In Table 5, the primary focus is on Formatting Quality, while Precision metrics serve as supplementary references. The Generalization metrics indicate the extent of catastrophic forgetting. When compared to the fine-tuned baseline PALR, $Ours_{full}$ shows the least performance deterioration from its underlying Llama2-7b architecture. Owing to the 512-token input limitation of InstructRec, its performance could not be assessed on the Generalization tasks.
3.4 Case Study
Instruction $\mathbf{I_{2}^{CP\approx 50\%}}$: You are an expert recommender engine. You need to generate a recommendation list simultaneously considering user‘s preference and user’s intention. Here are the user’s historical interactions: ’The Blind Swordsman: Zatoichi Takeshi Kitano; Tadanobu Asano japan’, ’Buffy the Vampire Slayer - The Complete Seventh Season’, ’Number Ones’, ’24: Season 3’, ’Madonna - The Immaculate Collection’, ’Spider-Man 2’, ’Heavy Metal 2000 VHS’, ’MI-5, Vol. 1’, ’MI-5, Vol. 2’, ’Beyond Re Animator VHS’, and user’s intention: In the recommendation list, the proportion of ’Art House & International, By Original Language, Chinese’ items should be approximately 50%. You need to generate a recommendation list with 5 different items.
$\mathbf{GPT-3.5}$ 1. Oldboy 2. Chungking Express 3. Battle Royale 4. In the Mood for Love 5. Memories of Murder
$\mathbf{Llama2-7b}$ 1. The Blind Swordsman: Zatoichi 2. Buffy the Vampire Slayer 3. Number Ones 4. 24: Season 3 5. Madonna - The Immaculate Collection
$\mathbf{InstructRec}$ 1. The Magnificent Seven VHS 2. Shanghai Noon 3. X-Men: First Class 4. X-Men VHS 5. Akira VHS
$\mathbf{Ours_{full}}$ 1. Five Deadly Venoms VHS 2. Ip Man 2 3. The Village: Vista Series 4. Dawn of the Dead 5. Fong Sai Yuk VHS
Instruction $\mathbf{I_{1}^{-C}}$: You are an expert recommender engine. You need to generate a recommendation list simultaneously considering user‘s preference and user’s intention. Here are the user’s historical interactions: ’Yet Another Zombie Defense HD’, ’Battlefield™ V’, ’Tom Clancy’s Ghost Recon® Wildlands’, ’Agents of Mayhem’, ’Beholder’, ’HELLDIVERS™ Dive Harder Edition’, ’SUPERHOT’, ’Watch_Dogs® 2’, ’Deus Ex: Mankind Divided’, ’POSTAL 2’, and user’s intention: Don’t recommend me any ’Shooter’ products. You need to generate a recommendation list with 5 different items.
$\mathbf{GPT-3.5}$ 1. Stardew Valley 2. Hollow Knight 3. Cities: Skylines 4. Divinity: Original Sin 2 5. The Witness
$\mathbf{Llama2-7b}$ 1. Assassin’s Creed Odyssey 2. The Witcher 3: Wild Hunt 3. Horizon Zero Dawn 4. Far Cry 5 5. Life is Strange 2
$\mathbf{InstructRec}$ 1. POSTAL 2 2. POSTAL Redux 3. Half-Life 2 4. Resident Evil 0 5. POSTAL
$\mathbf{Ours_{full}}$ 1. The Witcher® 3: Wild Hunt 2. Human: Fall Flat 3. The Forest 4. Outlast 5. Middle-earth™: Shadow of War™
Table 6: Two real examples for case study. We highlight some texts for Repeat with history item, In target category item and target item.
For clarity, we present two illustrative cases. The first is about the $I_{2}^{CP\approx 50\%}$ instruction on the Movie dataset. A user requests that roughly 50% of the top-5 recommendations feature “Art House & International, By Original Language, Chinese” characteristics. Our method successfully generates a list where 3 out of 5 items possess these attributes, while three baseline methods fail to meet this criterion. Llama-2-7b simply repeats items mentioned in the user history.
In the second case, concerning the $I_{1}^{-C}$ instruction, the user specifies to exclude “Shooter” games from recommendations. Our method effectively adheres to this restriction by omitting any items from the blacklisted category and successfully includes the ground-truth item in the recommendation list. Conversely, Llama-2-7b incorrectly suggests “Far Cry 5”, a Shooter game, while InstructRec not only makes aching historical mistakes but also recommends a lot of games from the shooter genre.
4 Related Work
In recent years, the remarkable natural language processing capabilities of LLMs have inspired researchers to leverage them for recommendation tasks Fan et al. (2023). Early implementations utilized language models as feature extractors, creating knowledge-aware recommendation embeddings exemplified by models such as U-BERT Qiu et al. (2021) and UserBERT Wu et al. (2022). With the advent of generative models like GPT, the focus has shifted towards generative recommendation models that frame recommendations as natural language generation tasks Wu et al. (2023).
Initially, adaptation of LLMs to recommendation scenarios relied heavily on techniques like prompt engineering Gao et al. (2023); Sun et al. (2023) and contextual learning Dai et al. (2023); Liu et al. (2023). However, these LLMs often underperform in comparison to traditional recommendation models trained on task-specific data, prompting efforts to fine-tune LLMs for better alignment with recommendation tasks. P5 Geng et al. (2022) introduced a unified framework that integrates five recommendation tasks through fine-tuning on the FLAN-T5 model Raffel et al. (2020). Subsequently, InstructRec Zhang et al. (2023) tailored FLAN-T5 for various downstream recommendation tasks using instruction tuning. TALLRec Bao et al. (2023) fine-tunes the LLaMA model for binary recommendation tasks with limited training samples, while PALR Chen (2023) employs two types of instructions for instruction tuning to facilitate list-level recommendation generation.
Despite these advancements, current studies have not fully explored the potential of LLMs to enhance the interactivity of recommender systems. We aim to harness the instruction-following prowess of LLMs for recommendation tasks through fine-tuning, to create a conversational, controllable, and interactive recommender agent.
5 Conclusion
In conclusion, we introduce a novel alignment methodology for LLMs in recommender systems, which significantly enhances their ability to follow user instructions and reduces formatting errors. By incorporating a supervised learning stage with a novel suite of control tasks and a reinforcement learning stage tailored with custom reward signals, our approach effectively bridges the gap between LLM capabilities and the nuanced requirements of recommendation tasks. Experiments confirm that our method not only outperforms existing LLM-based recommender systems but also demonstrates promising capabilities towards controllable and accurate recommendation experiences.
6 Limitation
This paper’s primary constraints are as follows: (1) The emphasis is placed on improving the LLM’s ability to follow recommendation-related instructions. This focus may inadvertently compromise the LLM’s general capabilities, as revealed in Table 5. (2) In real-world scenarios, users often have diverse control intentions, including intricate blends of various instructions. As foundational research, this paper addresses only the most critical elements, specifically category control and formatting control.
Appendix A Appendix
A.1 Dataset detail
We present detailed information about the instruction set we constructed in Table 7, including the quantity of each type of instruction.
Instruct type Movie Steam
$I_{0}$ $13218$ $12658$
$(I_{1}^{+C},I_{1}^{-C})$ $(6609,6609)$ $(6329,6329)$
$(I_{2}^{CP\leq*},I_{2}^{CP\geq*},I_{2}^{CP\approx*})$ $(4406,4406,4406)$ $(4219,4219,4220)$
$I_{3}$ $18744$ $8572$
$I_{4}$ $29199$ $23273$
$ALL$ $87597$ $69819$
Table 7: Constructed instruction set
input : $items=[item_{1},...,item_{N}],k,C_{target}$
output : Each item’s score $Scores^{ctl}$
init : $Scores^{ctl}=\mathbf{0}^{1\times N}$
init : $Count_{in}=Count_{out}=0$
1 for $i\leftarrow 1$ to $N$ do
2       if $item_{i}$ $\mathrm{is}$ $\mathrm{illegal}$ then
3             $Scores^{ctl}_{i}=-1$
4             continue
5      if $item_{i}\in C_{target}$ then $in=1,out=0$;
6       else $in=0,out=1$;
7       $Count_{in}+=in,Count_{out}+=out$ if $I_{0}$ then $s=Scores_{i}$;
8       if $I_{1}^{+C}$ then $s=in$;
9       if $I_{1}^{-C}$ then $s=out$;
10       if $I_{2}^{CP\leq m}$ then
11             if $Count_{out}>(k-k*m)$ then $s=0.5$;
12             else if $out$ then $s=1.0$;
13             else if $Count_{in}<k*m$ then $s=0.5$;
14             else $s=0.0$;
15
16      if $I_{2}^{CP\geq m}$ then
17             if $Count_{in}>k$ then $s=0.5$;
18             else if $in$ then $s=1.0$;
19             else if $Count_{out}<(k-k*m)$ then $s=0.5$;
20             else $s=0.0$;
21
22      if $I_{2}^{CP\approx m}$ then
23             if $in$ then
24                   if $Count_{in}\leq k*m$ then $s=1.0$;
25                   else $s=0.0$;
26
27            else if $Count_{in}\geq k*m$ then $s=1.0$;
28             else if $Count_{out}\leq(k-k*m)$ then $s=0.5$;
29             else $s=0.0$;
30
31      $Scores^{ctl}_{i}=s$
Algorithm 1 Calculation of $Scores^{ctl}$
A.2 Prompts
In this section, we show the prompt for the instruction used in the experiment. The prompt of $I_{0},I_{1},I_{2},I_{3}$ is illustrated in 1 to 4 respectively. The prompt of Positive and Negative category control intentions is illustrated in 5 to 6 respectively
Listing 1: Prompts of $I_{0}$
Instruction: You are an expert recommender engine. You need to generate a recommendation list considering users preference from historical interactions. The historical interactions are provided as follows: {history}. You need to generate a recommendation list with {item_count} different items.
Output: {item_list}
Instruction: You are an expert recommender engine. You need to select a recommendation list considering users preference from historical interactions. The historical interactions are provided as follows: {history}. The candidate items are: {candidate_titles}. You need to select a recommendation list with {item_count} different items from candidate items.
Output: {item_list}
Listing 2: Prompts of $I_{1}$
Instruction: You are an expert recommender engine. You need to generate a recommendation list simultaneously considering users preference inferred from historical interactions and users intention. If users preference conflicts with his intention, you should comply with his intention. Here are users historical interactions: {history}, and users intention: {synthetic_intention}. You need to generate a recommendation list with {item_count} different items.
Output: {item_list}
Instruction: You are an expert recommender engine. You need to select a recommendation list from candidate items simultaneously considering users preference inferred from historical interactions and users intention. If users preference conflicts with his intention, you should comply with his intention. Here are users historical interactions: {history}, and users intention: {synthetic_intention}. The candidate items are: {candidate_titles}. You need to select a recommendation list with {item_count} different items from candidate items.
Output: {item_list}
Listing 3: Prompts of $I_{2}$
Instruction: You are an expert recommender engine. You need to generate a recommendation list simultaneously considering users preference inferred from historical interactions and users intention. Here are users historical interactions: {history}, and users intention: In the recommendation list, the proportion of ’{target_category}’ items should be less than or equal to {category_proportion}. You need to generate a recommendation list with {item_count} different items.
Output: {item_list}
Instruction: You are an expert recommender engine. You need to generate a recommendation list simultaneously considering users preference inferred from historical interactions and users intention. Here are users historical interactions: {history}, and users intention: In the recommendation list, the proportion of ’{target_category}’ items should be more than or equal to {category_proportion}. You need to generate a recommendation list with {item_count} different items.
Instruction: You are an expert recommender engine. You need to generate a recommendation list simultaneously considering users preference inferred from historical interactions and users intention. Here are users historical interactions: {history}, and users intention: In the recommendation list, the proportion of ’{target_category}’ items should be approximately {category_proportion}. You need to generate a recommendation list with {item_count} different items.
Output: {item_list}
Listing 4: Prompts of $I_{3}$
You are an expert recommender engine. You need to generate a recommendation list complying users intention. Here is users intention: {synthetic_intention}. Please generate a recommendation list with {item_count} different items.
Output: {item_list}
You are an expert recommender engine. You need to select a recommendation list complying users intention from candidate items. Here is users intention: {synthetic_intention}. The candidate items are: {candidate_titles}. Please select a recommendation list with {item_count} different items from candidate items.
Output: {item_list}
Listing 5: Prompts of positive intention in $I_{1}$ and $I_{3}$
I like ’{target_category}’ products
Please recommend some ’{target_category}’ items
Im interested in ’{target_category}’
I would like to buy some ’{target_category}’ products
I would like to browse some ’{target_category}’ products
I prefer in ’{target_category}’ item
Listing 6: Prompts of negative intention in $I_{1}$ and $I_{3}$
I dont like ’{target_category}’ products
Please exclude any ’{target_category}’ item
Im not interested in ’{target_category}’
Dont recommend me any ’{target_category}’ products
I dont want to browse any ’{target_category}’ product
I hate ’{target_category}’ items